A clothing manufacturer has some pure silk thread and some thread that is 85% silk. How many kilograms of each must be woven together to make 75kg. of a blend that is 96% silk?
Answer by stanbon(48551)

A clothing manufacturer has some pure silk thread and some thread that is 85% silk. How many kilograms of each must be woven together to make 75kg. of a blend that is 96% silk?
-------------
85% thread DATA:
Amt. of thread = x kg ; Amt of silk in it= 0.85x kg
------------------
Pure silk DATA:
Amt of thread = "75-x"kg ; Amt of silk in it = 1(75-x)kg
---------------------
Mixture DATA:
Amt of mixture = 75 kg ; Amt of silk in it = 0.96*75 = 72 kg
-----------------------
EQUATION:
silk + silk = silk
0.85x + 75-x = 72
-0.15x = -3
x = 20 kg (Amt of 85% thread in the mixture)
------
75-x = 55 kg (Amt of pure silk in the mixture)
===================
Cheers,
Stan H.